ο‚· Demand:𝒑𝒅 =πŸ•πŸβˆ’πŸ.πŸπŸ“βˆ—π’’π’…

ο‚· Supply:𝒑𝒔 =𝟏𝟐.πŸ“+𝟐.πŸπŸ“βˆ—π’’π’”

Imagine now that the government imposes a tax of $πŸ’. πŸ“ on the producers of hammers.

a) What is the new equilibrium quantity?


1
Expert's answer
2022-03-13T19:02:34-0400

At equilibrium, "p_{d}" ="q_{s}" and "q_{d}" ="q_{s}" , demand= supply

71 - 1.25"q_{d}" = 12.5 + 2.25"q_{s}"

"\frac{3.5q}{3.5}" = "\frac{59.5}{3.5}"

q = 17 which is the equilibrium quantity

"p_{d}" = 72 - ( "1.25\times17)"

"p_{d}" =$ 50.75 (equilibrium price)

Imposition of tax will increase the equilibrium price and thereafter the new price will be;

p= $50.75+$4.5 =$ 55.25

the demand function is,"p_{d}" = 72-1.25"q_{d}"

Replacing pd with the new price we have,

55.25 = 72 - 1.25 "q_{d}"

"q_{d}" = $13.4(new equilibrium quantity)